Given an actual demand of 107, a previous forecast value of 109, and an alpha of 0.82, the exponential smoothing forecast for the next period would be (rounded to 2 decimals)
Answer #1

For this the formula is as follows:

Ft+1=αAt+ (1-α) Ft

Here α=0.82

At=Actual Demand=107

Ft= Previous forecast value=109

Ft+1       = (0.82) (107) + (1-0.82) 109

            = 87.74+ (0.18)109

            =87.74+19.62

            =107.36

The required answer is 107.36